OpenStudy (anonymous):

a cone that holds slushy ice in the snow has a diameter of 6cm and a height of 7cm. if the ice that fills the CONE rises above the cone to form a perfect hemisphere, calculate the volume of the slushy ice? PLEASE HELP. :( last question for a sunday night hw

OpenStudy (campbell_st):

the radius of the cone is r = 3 cm and the height is h = 7 cm the volume of a cone is \[V = \frac{1}{3} \pi r^2 h\] the volume of the hemisphere is \[V = \frac{1}{2} \times \frac{4}{3} \pi r^3\] add the 2 volumes together to get the volume of ice